Prove that the values of the polynomial x^3–x are divisible by 6 for whole values of x.
Answer by ikleyn(52879)

.
x^3 - x = x*(x-1)*(x+1),


so the left side is the product of three consecutive integer numbers.


Of them, at least one is a multiple of 2 and exactly one is a multiple of 3.


Hence, the product, which is x^3 - x, is divisible by 2*3 = 6.

Answered and solved.